OpenStudy (iwanttogotostanford):

HELP ASAP WITH THIS QUESTION:

OpenStudy (iwanttogotostanford):

@Kevin @mathmate @welshfella

OpenStudy (3mar):

How is it solved @Kevin?

OpenStudy (welshfella):

divide top and bottom of the fraction by x^4

OpenStudy (3mar):

then?

OpenStudy (welshfella):

then you'll have 2 fractions to deal with

OpenStudy (kevin):

You cannot factorize or find its derivative to get a certain number. So what you do is just fill x with 0 \[\frac{ x^2 + 3 }{ x^4 }\] \[\frac{ 0^2 + 3 }{ 0^4 }\] \[\frac{ 3 }{ 0 }\]

OpenStudy (3mar):

Try what welshfella siad "divide top and bottom of the fraction by x^4"

OpenStudy (3mar):

It does make sense.

OpenStudy (iwanttogotostanford):

so the limit would be 3???

OpenStudy (kevin):

3/0 = infinity

OpenStudy (iwanttogotostanford):

oh so the limit if infinity?? Or how d o you word that?

OpenStudy (kevin):

you know infinity symbol right? 

OpenStudy (iwanttogotostanford):

yes

OpenStudy (kevin):

that's the answer

OpenStudy (kevin):

I'm not good actually. lol